OpenStudy (anonymous):

2t^3/4t+8 times t+2/8t

OpenStudy (anonymous):

Hello are you there

OpenStudy (anonymous):

Solve for t over the real numbers: 5/8 t (t^4+16) = 0 Multiply both sides by a constant to simplify the equation. Multiply both sides by 8/5: t (t^4+16) = 0 Solve each term in the product separately. Split into two equations: t = 0 or t^4+16 = 0 Look at the second equation: Isolate terms with t to the left hand side. Subtract 16 from both sides: t = 0 or t^4 = -16 Show the equation has no solution. t^4 = -16 has no solution since for all t on the real line, t^4 = t^4 >=0 and -16ɘ: Answer: | | t = 0